a mother was 26 years old when her daughter was born. Today the sum of three times her age plus twice her daughters age is 103 years. how old are they today
Answer by ankor@dixie-net.com(22740)

Let m = Mom's present age
Let d = daughter's
:
Write an equation for each statement
:
"a mother was 26 years old when her daughter was born."
This means
m = d + 26
:
"Today the sum of three times her age plus twice her daughters age is 103 years."
3m + 2d = 103
Replace m with (d+26)
3(d+26) + 2d = 103
3d + 78 + 2d = 103
5d = 103 - 78
5d = 25
d = 25/5
d = 5 yrs is daughter's age
and
m = 5 + 26
m = 31 yrs is Mom's age
:
:
Check solutions in the statement:
"The sum of three times her age plus twice her daughters age is 103 years."
3(31) + 2(5) = 103